# Defect configuration of an active nematic around a circular obstacle

### v1

@@ Abstract
Nematic liquid crystal is a complex fluid whose constituents have an anisotropic shape and exhibit orientational order. A class of active matter which exhibits orientational order, e.g., cell populations, is called active nematic. A singularity of the orientational field, called a topological defect, is a robust structure which cannot be removed by continuum deformations. This robustness plays an important role in, e.g., the process of morphogenesis. To enhance the understanding of the behavior of active nematic, it is important to understand the behavior of topological defects. In this paper, we study the configuration of topological defects of a two-dimensional active nematic around a circular obstacle. In the case of a passive nematic liquid crystal, the equilibrium configuration of defects can be easily identified by the method of image charges. In the case of an active nematic, however, one must take account of the flow field generated by active constituents, and the problem of identifying the defect configuration becomes complicated. In this study, we perform numerical simulations and investigate how the stationary defect configuration deviates from the passive case. Furthermore, we carry out a theoretical calculation based on an analytical expression relating the defect velocity with the force exerted on the defect. Our theoretical calculation qualitatively reproduces the simulation results. Our study may be applied to describing the behaviour of e.g. cell populations in the presence of obstacles, and has the potential to benefit related fields, e.g., developmental biology.
@@ II.1 Basic equations
The first term ($\mu\bm{u}$) in the r.h.s. of eq.(2) is the friction between active nematic and the substrate. The second term ($\eta\bm{\nabla}^{2}\bm{u}$) is the shear viscosity, and the third term ($\alpha\bm{\nabla}\cdot\bm{Q}$) represents the active stress. The magnitude of $\alpha$ represents the strength of the activity. $\alpha<0$ and $\alpha>0$ correspond to the extensile and contractile active nematic, respectively [2]  [^2] We note the difference between the definitions of the activity parameter in [2] and this paper. In [2], the active stress tensor is defined as $\zeta\bm{Q}$ and thus $\zeta>0$ and $\zeta<0$ describe the extensile and contractile active nematic, respectively . As in [13, 14, 9], we use the compressible but constant density limit, where the density is assumed to be constant, but $\bm{\nabla}\cdot\bm{u}=0$ is not enforced. This means that the mass conservation is broken but such a situation may be realized by allowing the birth and death processes of constituents. The pressure is assumed to depend only on the density, which is assumed to be constant, and does not appear in the Stokes equation (Eq.(2)). Strictly, the bulk viscosity term ($\propto\bm{\nabla}\left(\bm{\nabla}\cdot\bm{u}\right)$) should be included in the Stokes equation (Eq.(2)) since we do not enforce $\bm{\nabla}\cdot\bm{u}=0$. However, this term will make our analysis of the Stokes equation (see Sec.III.3.2 or Appendix D) quite complicated. Thus, we ignore this term in this paper.

### v2

@@ Abstract
To enhance the understanding of the behavior of active nematic, it is important to understand the behavior of topological defects. In this paper, we study the configuration of topological defects of a two-dimensional active nematic around a circular obstacle. In the case of a passive nematic liquid crystal, the equilibrium configuration of defects can be easily identified by the method of image charges. In the case of an active nematic, however, one must take account of the flow field generated by active constituents, and the problem of identifying the defect configuration becomes complicated. We first perform numerical simulations and investigate how the stationary defect configuration deviates from the passive case. Furthermore, we carry out a theoretical calculation based on an analytical expression relating the defect velocity with the force exerted on the defect. Our theoretical calculation qualitatively reproduces the simulation results. Our study may be applied to describing the behaviour of e.g. cell populations in the presence of obstacles, and has the potential to benefit related fields, e.g., developmental biology.
@@ II.1 Basic equations
The first term ($\mu\bm{u}$) in the r.h.s. of eq.(2) is the friction between active nematic and the substrate. The second term ($\eta\bm{\nabla}^{2}\bm{u}$) is the shear viscosity, and the third term ($\alpha\bm{\nabla}\cdot\bm{Q}$) represents the active stress. The magnitude of $\alpha$ represents the strength of the activity. $\alpha<0$ and $\alpha>0$ correspond to the extensile and contractile active nematic, respectively [2]  [^2] We note the difference between the definitions of the activity parameter in [2] and this paper. In [2], the active stress tensor is defined as $\zeta\bm{Q}$ and thus $\zeta>0$ and $\zeta<0$ describe the extensile and contractile active nematic, respectively . As in [13, 14, 9], we use the compressible but constant density limit, where the density is assumed to be constant, but $\bm{\nabla}\cdot\bm{u}=0$ is not enforced. This means that the mass conservation is broken but such a situation may be potentially realized by allowing the birth and death processes of constituents. The pressure is assumed to depend only on the density, which is assumed to be constant, and does not appear in the Stokes equation (Eq.(2)). The simulations under the above-mentioned compressible but constant density limit give qualitatively similar results to the ones obtained in particle-based simulations of dry truly compressible fluids [13]. Furthermore, strictly, the bulk viscosity term ($\propto\bm{\nabla}\left(\bm{\nabla}\cdot\bm{u}\right)$) should be included in the Stokes equation (Eq.(2)) since we do not enforce $\bm{\nabla}\cdot\bm{u}=0$. However, this term will make our analysis of the Stokes equation (see Sec.III.3.2 or Appendix D) quite complicated. Thus, we ignore this term in this paper.
@@ III.3.4 Comparison between simulation results and theoretical calculations
Furthermore, let us discuss the point that the agreement between theory and simulation is not improved when the higher-order terms are included; see $\alpha>0$ regime of FIG.7, where the theory and simulation show best agreement when the infinite summation in Eq.(16) is truncated up to $n=1$. Although one might expect that the theoretical curve gets closer to the simulation results when the higher-order ($n=2,3,\cdots$) terms are included, it is not the case. Our theoretical framework is based on several approximations listed in Sec.III.3.2. If the difference between the theory and simulation arising from those approximations decreased by including the higher-order terms, the theoretical curve would get closer to the simulation results. However, it is not the case in our framework, and we cannot expect that the agreement increases. Our result of homeotropic case (FIG.7) indicates that the approximations do not improve when higher-order terms are included. We would like to leave further elaboration of our theory for future studies.